Y=√x cos 2x find dy/dx

1
Expert's answer
2022-06-02T10:26:00-0400
"y=\sqrt{x}\cos2x"

"\frac{dy}{dx}=\frac{1}{2\sqrt{x}}\cos2x-2\sqrt{x}\sin2x"